OpenStudy (anonymous):

5 + 6x - 3y + x + 8y

OpenStudy (anonymous):

?? 5+7x-5y i didnt fully understand the question??

OpenStudy (anonymous):

sorry what is it in simplest form

OpenStudy (anonymous):

Add '-5' to each side of the equation 7x + 5y = -5 Add '-5y' to each side of the equation. 7x = -5 + -5y Divide each side by '7' i think thats it

OpenStudy (anonymous):

5 + 6x - 3y + x + 8y = 5+7x+5y

OpenStudy (anonymous):

-4x - 4 + x - 2y + 7x here x has three terms, add them togetehr

OpenStudy (anonymous):

-4x - 4 + x - 2y + 7x = -4x+x+7x-2y-4 =(-4+1+7)x -2y -4